OpenStudy (anonymous):

A system of equations is shown below: x + 3y = 5 (equation 1) 7x – 8y = 6 (equation 2) A student wants to prove that if equation 2 is kept unchanged and equation 1 is replaced with the sum of equation 1 and a multiple of equation 2, the solution to the new system of equations is the same as the solution to the original system of equations. If equation 2 is multiplied by 1, which of the following steps should the student use for the proof?

OpenStudy (anonymous):

(A) Show that the solution to the system of equations 3x + y = 5 and 8x –7y = 6 is the same as the solution to the given system of equations (B) Show that the solution to the system of equations 8x – 5y = 11 and 7x – 8y = 6 is the same as the solution to the given system of equations (C) Show that the solution to the system of equations 15x + 13y = 17 and 7x – 8y = 6 is the same as the solution to the given system of equations (D) Show that the solution to the system of equations –13x + 15y = 17 and 7x – 8y = 6 is the same as the solution to the given system of equations

OpenStudy (anonymous):

@paki I think the answer is either B or D. Can you confirm plz?

OpenStudy (anonymous):

@dan815 @TheSmartOne

OpenStudy (nincompoop):

THINK? HOW ABOUT YOU SOLVE IT

OpenStudy (anonymous):

I did solve it on paper and pencil

OpenStudy (nincompoop):

show your solution

OpenStudy (anonymous):

Here let me show you why

OpenStudy (anonymous):

Because the directions said "equation 2 is kept unchanged and equation 1 is replaced with the sum of equation 1 and a multiple of equation 2," it means that you add x + 3y = 5 (equation 1) and 7x – 8y = 6 (equation 2) together. You will get 8x -5y = 11

OpenStudy (anonymous):

And then the directions say, "equation 2 is multiplied by 1." Multiplying equation 2 by 1 will leave it the same: 7x – 8y = 6

OpenStudy (nincompoop):

what is equation 1 is replaced with sum of equation 1?

OpenStudy (anonymous):

^equation 1 is replaced with the sum of equation 1 and a multiple of equation 2

OpenStudy (anonymous):

will be 8x -5y = 11

OpenStudy (nincompoop):

can you rewrite that in equation

OpenStudy (nincompoop):

where is the "multiple"

OpenStudy (anonymous):

The multiple of equation 2, I am assuming, is the same as equation 2: 7x – 8y = 6

OpenStudy (anonymous):

x + 3y = 5 7x – 8y = 6 ----------- 8x - 5y = 11